OpenStudy (anonymous):

a 2000lb car is on a hydraulic lift, what it the pressure is the side of the lift has a 20cm diameter?

OpenStudy (anonymous):

Pressure = Force / Area

OpenStudy (anonymous):

convert units to SI to make it easier

OpenStudy (anonymous):

2000lb (kg ) = 909kg 2.2lb

OpenStudy (anonymous):

Area... d = 20 cm = .2m A = pi r^2 = pi .1m^2 = .0314m^2

OpenStudy (anonymous):

Force = mass x gravity = 909kg x 10m/s^2 = 9090N

OpenStudy (anonymous):

P = F/A P = 9090N/.0314m^2 = 289,490Pa

OpenStudy (anonymous):

There's your answer :)

OpenStudy (anonymous):

Medal?